Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2020-01-17

Observers: Erez Zimmerman, Ido Irani, Steve Schulze
Support Team: Rachel Bruch, Ofer Yaron
TAT: Nancy Elias-Rosa, Giuliano Pignata, Morgan Fraser

Calibrations: 

Weather: 
Start: Humidity 30%, wind speed 1.7 m/s, seeing: 1".
Middle: Humidity 20%, wind speed 5 m/s, seeing: 1.22".
End : Humidity 23%, wind speed 2 m/s, seeing: 1.22".

Comments:
Problem with the connection between the VOT and P2.
The compressor of the cooling unit of SOFI broke in the middle of the night and SOFI started to warm up. The warming-up prevented us from taking a telluric.
End of the second half is affected by clouds.

Observations:


UT Target Type Setup Seeing Exptime Notes
01:02GD71STD
1.0

01:30AT2019qemFollow-upJHK0.831480 s
AT2019xisFollow-upJHK0.833960 s
03:05AT2019esaFollow-upBG+RG0.585760 stechnical problems, aborted
05:06Problem with SOFI
compressor failed
05:12HIP 40415Telluricaborted b/c SOFI is warming up
05:22Switch to EFOSC2
05:42LTT3864STDgr11+gr13+gr161.22
06:10AT2019unbFollow-upgr131.222x1800 s
07:30SN2019wbgFollow-upgr131.221800 sclouds
08:03SN2020oiFollow-upgr11+gr161.22abortedclouds, going to the next target
08:11AT2019zcrClassificationgr131.22abortedclouds







Comments